Exercise 15-4 Riverbed Corporation is a regional company which is an SEC registrant. The corporations securities are thinly traded on NASDAQ. Riverbed Corp. has issued 16,000 units. Each unit consists of a $800 par, 12% subordinated debenture and 16 shares of $8 par common stock. The units were sold to outside investors for cash at $1,408 per unit. Prior to this sale, the 2-week ask price of common stock was $64 per share. Twelve percent is a reasonable market yield for the debentures, and therefore the par value of the bonds is equal to the fair value. (a) Prepare the journal entry to record Riverbeds transaction, under the following conditions. (Round answers to O decimal places, e.g. $38,487. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select No Entry for the account titles and enter O for the amounts.) (1) Employing the incremental method. (2) Employing the proportional method, assuming the recent price quote on the common stock reflects fair value. No. Account Titles and Explanation Debit Credit 1. 2.
Answer #1

Solution:

Journal entries to record Reverded's transaction.

No Account titles and explanation Debit($) Credit($)
1 Cash A/c Dr (16,000 * 1,408) 22,528,000
To Bonds payable A/c (16,000 * 800) 12,800,000
To Common stock A/c (16,000 * 16 * 8) 2,048,000
To Paid in capital excess of par- common stock A/c 7,680,000
2 Cash A/c Dr 22,528,000
Bonds discount A/c Dr [(16,000*800) - 9,880,780.8] 2,919,219
To Bonds payable A/c 12,800,000
To Common stock A/c 2,048,000
To Paid in capital excess of par- common stock A/c (12,647,219 - 2,048,000) 10,599,219

Working notes:

Debentures = (16,000 * 1,408) * 800 / (800 + 16 * 64)

= 22,528,000 * 800 / 1,824

= 22,528,000 * 0.4386

= 9,880,780.8

Common shares = (16,000 * 1,408) * (1,024 / 1,824)

= 22,528,000 * 0.5614

= 12,647,219